OpenStudy (anonymous):

hi can u help me, please A 2Kg object is moving at 5m/s. A 8N force is applied in the direction of motion and then removed after the object has traveled an additional 2s. Calculate the work done on it.

OpenStudy (anonymous):

The answer is 144 J . How can I get it ????

OpenStudy (anonymous):

The Work equal \[W(F)_{a-b}=F.d.\cos \alpha \] F is the force (N) d is the distance cos alpha if Vertically you dont use the cos alpha

OpenStudy (anonymous):

This formula is used only when the force is constant !

OpenStudy (anonymous):

of course but i dont know why your force is not constant

OpenStudy (anonymous):

i think the key to solve this is to consider that \[W = \Delta K = \frac{ 1 }{ 2 } m (v_{f}^{2} - v_{i}^{2} )\] but i dont know how can i get the final velocity ?!!!

OpenStudy (anonymous):

I'm sorry I'm not at this level ...

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

u can do this in ur method using work-energy theorem or using method by @AntarAzri if u do by ur method,find the final velocity for the time t=2s using v=u+at or by the other: u should not that work is done only when a force acts onb a body so initially the body moved with CONSTANT velocity=== no force so find the distance it moves in the presence of a force s=ut+1/2at^2 then done